# Analyzing GitHub Issues and Pull Requests in nf-core Pipelines: Insights into nf-core Pipeline Repositories

### v1

@@ 3.2. Pre-processing Issues and Pull Requests
To enable reliable topic modeling of GitHub data, we perform a systematic preprocessing pipeline to clean and normalize the textual data. Specifically, we concatenate the title and body of each issue or PR into a single document to preserve sufficient contextual information for topic extraction. We then remove code snippets enclosed in backticks using regular expressions, as such blocks often contain syntax elements irrelevant to semantic modeling. Similarly, we eliminate URLs, which typically link to external resources, and clean out residual HTML tags. We further normalize the corpus by removing special characters and punctuation. To address high-frequency but semantically uninformative words, we apply stopword removal using the NLTK stopword list (Hardeniya et al., 2016), and supplement it by filtering out project-specific terms such as nf-core and nextflow to prevent these frequently occurring system names from dominating the topic space. Finally, we apply lemmatization (Balakrishnan and Lloyd-Yemoh, 2014) using the spaCy en_core_web_sm model (Vasiliev, 2020) to convert words into their base forms, thus improving topic coherence and reducing lexical variation. This preprocessing ensures that the data retains its semantic richness while reducing noise, making it suitable for subsequent topic modeling and analysis. After preprocessing, we identify three samples with empty titles and bodies (these originally contained only URLs in the title and empty body). As they provide no meaningful information, we exclude them from the analysis. Consequently, the final dataset consists of 25,173 issues and PRs.
@@ 3.3. Identifying Topics in nf-core Pipelines
Determining topic modeling parameters: In BERTopic, the nr_topics parameter allows users to control the number of topics by merging similar ones after their initial creation. However, prior work (Alam et al., 2025; Li et al., 2021) has emphasized the value of relying on clustering algorithms to produce a more natural topic structure. To enhance clustering quality, we first reduce the dimensionality of the document embeddings, as high-dimensional data can hinder performance. For this step, we employ UMAP (McInnes et al., 2018), which effectively preserves local semantic structure while projecting embeddings into a lower-dimensional space. We then apply HDBSCAN (McInnes et al., 2017), a density-based algorithm particularly well suited to UMAP projections. Unlike traditional clustering methods, HDBSCAN is capable of identifying noise points and does not force every document into a cluster, making it especially effective for GitHub data.
For BERTopic modeling, we systematically tune UMAP and HDBSCAN parameters to balance topic granularity with semantic coherence. To evaluate the resulting topics, we employ the topic coherence score (Röder et al., 2015), a widely used metric that measures the semantic similarity among the top words in a topic. This measure helps distinguish between topics that are semantically interpretable and those that are merely artifacts of statistical inference. Specifically, for UMAP we vary n_neighbors (15–50) to balance local semantic similarity with a meaningful global structure, and adjust n_components (5–30) to retain sufficient semantic information while enabling efficient clustering. We use cosine distance, which aligns well with sentence-transformer embeddings. For HDBSCAN, we experiment with min_cluster_size values between 30 and 200, adopting Euclidean distance as it is well suited to UMAP-reduced spaces. To improve topic representation, we apply CountVectorizer with unigram and bigram models, consistent with prior studies (Alam et al., 2025; Li et al., 2021; Abdellatif et al., 2020).
We achieve the highest topic coherence score of 0.904 for our dataset using n_neighbors = 30, n_components = 15, and min_cluster_size = 150. With this configuration, BERTopic produces 13 distinct topics. However, BERTopic introduces a special -1 topic to aggregate documents that do not align well with any of the discovered topics. Such cases typically occur when documents contain excessive noise (e.g., stopwords or irrelevant content) or lack strong connections to the main themes extracted by the model. Consequently, the -1 topic acts as a catch-all cluster for outlier documents. Since our dataset underwent extensive preprocessing to remove noisy and irrelevant elements prior to topic modeling, our main topic began with the label -1. The preprocessing and topic generation scripts are available in our replication package (Random, 2025). A detailed analysis of the resulting topics is presented in the Section 4.
@@ 4. Results
In this section, we analyze the issues and PRs from the nf-core pipelines. Our experiments are designed to address the following three research questions:
  • RQ1: What topics do users raise and discuss in the issues and PRs of nf-core pipelines?
  • RQ2: How are issues and PRs managed and addressed in nf-core pipelines, and how do different features influence their resolution?
  • RQ3: To what extent do developers perceive the revealed challenges in terms of difficulties?
@@ 4.1. RQ1: What topics do users raise and discuss?
Motivation: Understanding the topics raised in issues and PRs is essential to identifying the challenges and needs in nf-core pipeline development and maintenance. Since these pipelines are community-driven and widely used in bioinformatics, the issues and PRs often reflect recurring technical difficulties, usability concerns, feature requests, and collaborative practices. Importantly, the development of nf-core pipelines differs substantially from that of traditional software systems: pipelines must integrate heterogeneous bioinformatics tools, manage complex data dependencies, and ensure reproducibility across diverse computational environments (e.g., containers, HPC clusters, cloud platforms). These requirements introduce domain-specific challenges, such as compatibility with reference datasets, versioned tools, and workflow execution engines, which are less prevalent in general-purpose software engineering (Di Tommaso et al., 2017; Arvanitou et al., 2021; Gruening et al., 2019). Therefore, the challenges that developers and users face in scientific workflow repositories like nf-core may differ from those commonly encountered in traditional software projects. The motivation behind this RQ1 is to reveal the challenges, needs, and recurring concerns of the nf-core pipeline repositories.
Approach: To identify the key topics, we apply BERTopic modeling to our datasets, as outlined in Section 3. The modeling process yields 13 distinct topics along with the distribution of co-occurring words for each. We then undertake a manual labeling procedure to assign meaningful and contextually accurate labels to the identified topics. Following established practices in prior work (Alam et al., 2025; Li et al., 2021; Openja et al., 2020; Yang et al., 2016; Bagherzadeh and Khatchadourian, 2019; Scoccia et al., 2021), the first author, having over five years of experience in nf-core pipelines and more than a decade of professional software development experience, proposes initial candidate labels. These are based on topic keywords and a manual review of at least 30 randomly selected representative issues and PRs per topic. The proposed labels are subsequently refined through collaborative discussions with the second author, who has over nine years of experience in bioinformatics-related pipeline research, and an additional expert with more than two decades of expertise in empirical software engineering research. Through this iterative process, we finalize a set of labels that accurately capture the underlying semantics and contextual nuances of the identified topics.
@@
1. Pipeline Development and Integration: This is the most dominant topic users face when working with nf-core pipelines, focusing on the foundational tasks of initiating and integrating new workflows in the nf-core ecosystem. For example, in the hlatyping pipeline, an issue, nf-core pipelines should provide structured output information, proposed extending nf-core metadata to not only describe inputs but also provide structured output information, helping humans and automated systems anticipate outputs, and route results. Developers frequently scaffold pipelines using standardized templates, add files and test stubs, and ensure compliance with nf-core community guidelines. Keywords such as new tool, add test, pipeline, and documentation update highlight a strong emphasis on code quality and adherence to standards, with many issues and PRs focusing on early development tasks like structuring projects.
2. Template Sync. and Automated Update Management: This is the second most frequent challenge, highlighting the need to keep pipelines synchronized with the latest nf-core template. Issues and PRs in this topic involve propagating new conventions and updates, resolving merge conflicts, and ensuring community standards. For instance, the dualrnaseq pipeline issue titled as Pipeline needs updating to the latest template version required fixing the TEMPLATE branch and creating a sync PR to resolve conflicts. Keywords such as conflict branch, make resolve, and automate attempt illustrate the hybrid nature of this process, where automation drives standardization but human intervention remains essential.
@@
5. Bug Fixing: Users often encounter challenges in ensuring pipeline reliability by resolving bugs, adding tests, and maintaining representative datasets (e.g., fix bug, add test, nfcoretestdataset). For example, the PR Bugfix MALT error retries addresses a memory-related issue by enabling the pipeline to retry when Java exits with code 1 due to heap space exhaustion, rather than aborting immediately.
6. Coordinating Contributions and Maintaining Pipeline Quality: Users often struggle to balance multiple tasks when contributing, such as updating documentation, fixing issues, and adding tests. Keywords like contribute, make branch, and appropriate delete reflect recurring coordination challenges across branches and the cleanup of outdated content. For example, the PR Revert multiqc workaround restores original functionality by removing a temporary fix after a patch release resolved the underlying issue. This topic illustrates the socio-technical complexity of nf-core development, where accuracy, quality, and collaboration must be maintained simultaneously.
@@
10. Tool Development and Repository Maintenance: Users frequently encounter challenges when creating new tools and maintaining the repositories. Keywords such as make and new tool reflect to the complexity of initiating tool development, while add test, branch pipeline, and appropriate delete indicate difficulties in ensuring correctness, managing branches, and removing outdated files. For example, the PR Add Kraken2 processes to build database and filter read illustrates these efforts. In short, this topic highlights challenges of pipeline repository maintenance, where new tool creation and repository evolution must be coordinated across contributors.
@@
RQ1 Summary RQ1 investigates the topics that users raise and discuss in the issues and PRs of nf-core pipelines. We identify 13 distinct topics through BERTopic modeling and manual validation. These topics encompass areas such as pipeline development and integration, template synchronization, debugging execution failures, tool and documentation maintenance, contribution coordination, genome data integration, module and parameter updates, subworkflow development, CI configuration, reporting/visualization, and version management. The taxonomy demonstrates that user discussions in nf-core repositories span the entire workflow lifecycle.
4.2. RQ2: Management and resolution of issues and PRs in nf-core pipelines.
@@
Motivation: In nf-core pipelines, GitHub issues and PRs serve as the primary channels for reporting bugs, requesting features, and contributing improvements. They reflect both the technical challenges users encounter and the collaborative processes used to resolve them. Understanding how these artifacts are managed is essential, as their resolution directly affects pipeline reliability, usability, and sustainability. As highlighted in RQ1, users raise issues and PRs for diverse reasons, and ignoring these challenges can prevent users from successfully using these pipelines. While some cases, such as installing required packages, can be easily addressed, others, like tool development, CI management, and debugging, require coordinated effort and management. While prior studies have explored issues and PRs management in general open-source projects, pipeline repositories like nf-core pose unique, domain-specific challenges that may shape how issues are prioritized, resolved, or neglected. Investigating how these processes are managed and how features such as labeling, assignment, or code snippets influence resolution can provide important insights into the socio-technical dynamics of collaborative pipeline maintenance. Thus, RQ2 seeks to uncover the practices, challenges, and strategies that sustain effective issue/PR management in nf-core, offering lessons for both the nf-core community and the broader ecosystem of reproducible pipeline development.
@@
Figure 2. The distribution of (a) closed issues and PRs rate, (b) ignored issues and PRs rate, and (c) addressing time.
@@
How are the issues and pull requests addressed? We examine how issues and PRs are managed in our dataset. These issues and PRs are categorized into two groups, open (unresolved) and closed (resolved). Among 25,173 issues and PRs, 22,500 (89.38%) are closed and 2,673 (10.62%) remained open, with an overall average of 201.38 entries per repository. Figure 2(a) illustrates the distribution of the Closed Issue and PR Rate (CIR) across repositories. Notably, 1.61% of repositories have achieved a CIR of 1.0, indicating that all reported issues and PRs have been resolved. Moreover, repositories with a CIR exceeding 0.9 exhibit an average of 259.13 issues and PRs, suggesting that repositories with higher closure rates also tend to manage a larger volume of entries. Overall, 48.39% of repositories demonstrate a CIR greater than 0.9, highlighting that nearly half of the repositories maintain a consistently high resolution rate.
@@
We find that 60.88% of closed issues and PRs are self-closed, meaning that the individuals who reported them were able to resolve their concerns, often through subsequent discussion. In addition, we identify another type of case referred to as ignored issues and PRs. These are cases that only the individuals who raised them are interested in, and receive no response from other users. These ignored cases account for 4.64 % of all issues and PRs. We also define the ignored issues and PRs rate as the ratio of ignored cases for each repository. This metric provides insight into the extent of active engagement by repository maintainers in addressing cases. The distribution of this rate across repositories is presented in Figure 2(b).
@@
We also analyze issues and PRs addressing time, defined as the duration between creation and closure. Among 22,500 closed cases, the mean addressing time is 47.51 days with a standard deviation of 141.67 days, while the median is 2.89 days, indicating most are resolved quickly. However, a subset remains unresolved for extended periods. The maximum addressing time observed exceeds four years, highlighting significant variability in resolution dynamics. Figure 2(c) illustrates the distribution of the addressing time.
@@
To examine factors influencing issue and PR resolution time, we analyze two repository-level features following prior work (Yang et al., 2023): (1) the number of contributors and (2) the total number of issues and PRs. We hypothesize that a larger contributor base may reduce addressing time. A negative Spearman correlation of –0.198 (p = 0.0283) between contributors and median closing time indicates that repositories with more contributors resolve issues slightly faster. Although this effect is statistically significant, the practical impact remains small. We further hypothesize that repositories with more issues and PRs would exhibit longer addressing times, as a larger workload might burden maintainers. Contrary to this expectation, the results reveal a Spearman correlation of –0.218 (p = 0.0157) between the number of issues and PRs and the median closing time. This negative correlation suggests that repositories handling a greater volume of issues/PRs actually resolve them somewhat faster, though again, the effect size is modest.
@@
How are the issues and pull requests managed? GitHub offers several mechanisms to support maintainers in managing issues and PRs. Following prior work (Yang et al., 2023), we focus on two commonly used strategies: (1) labeling, where issues and PRs are tagged with labels, and (2) assigning, where issues and PRs are allocated to specific individuals for resolution. In the remainder of this section, we use the terms labeling and assigning to refer to these two practices.
@@
We observe that 33.23% (8,366) of issues and PRs are tagged with labels, while 17.13% are assigned to specific contributors. This indicates that a large portion of issues and PRs do not make use of GitHub’s built-in management features. To further examine the use of labels, we note that GitHub provides nine default options: bug, documentation, duplicate, enhancement, help wanted, invalid, question, wontfix, and good first issue. Actively managed projects often go beyond these defaults by defining additional labels (e.g., DSL2). Across all pipeline repositories, we identify 154 distinct labels (including the defaults). We then calculate the frequency of each label, with the ranked distribution presented in Figure 3.
@@
Relationships between different features and the closure of issues and PRs: We further examine how different characteristics of issues and PRs relate to their likelihood of closure. Understanding these relationships offers practical insights for repository maintainers and contributors on improving the management and resolution of issues and PRs. Specifically, we focus on the following features:
@@
1. has-label: Indicates whether an issue/PR has a label. On GitHub, users with triage access can apply labels to categorize or prioritize Issues/PRs. A labeled entry usually means that it has been noticed and read by the maintainers, thereby increasing the likelihood of being actively addressed.
@@
2. has-assignee: Denotes whether an issue or PR is assigned to a contributor. On GitHub, users with write access can assign items to specific developers. An assigned issue/PR typically indicates clear responsibility, which may increase the likelihood of timely and successful resolution.
@@
3. title-length: Refers to the length of an issue or PR title. Informative, longer titles provide more context, helping maintainers understand and resolve items more quickly.
@@
4. body-length: Represents the issue or PR description length. Detailed descriptions give maintainers essential context, aiding triage and resolution, and are often encouraged by project guidelines.
@@
5. has-code: Indicates if the issue or PR includes code blocks, which help maintainers reproduce and resolve issues more effectively.
@@
6. has-URL: Indicates if the issue or PR contains URLs, which provide helpful context through links to related issues, documentation, or external resources.
@@
7. num-contributors: Refers to the total number of contributors in a repository. Intuitively, repositories with a larger contributor base have greater collective capacity, which may increase the chances of timely issue/PR resolution.
@@
Feature Measurement and Statistical Test: We compute the length of the issue/PR title and body by converting them into Python strings. Using regular expressions, we further identify whether an issue/PR description contains code blocks or URLs. For every issue/PR, we measure the features above mentioned. The number of contributors is obtained by querying the GitHub API to fetch metadata about the repository. Finally, we apply the Wilcoxon rank-sum test (Wilcoxon, 1945) to assess the statistical significance of the differences in feature distributions between open and closed issues and PRs.
@@
Table 2. The significance of the differences between the feature values of open and closed issues and PRs (Wilcoxon rank-sum test and Cohen’s effect sizes $\delta$).
Feature P-value Cohen’s $\delta$
@@
We also estimate Cohen’s effect size ($\delta$) to quantify the magnitude of differences between open and closed issues and PRs. Following established guidelines, values of $|\delta|<0.2$ are regarded as negligible, $0.2\leq|\delta|<0.5$ as small, $0.5\leq|\delta|<0.8$ as medium, and $|\delta|\geq 0.8$ as large. Table 2 reports the results of the Wilcoxon rank-sum test and corresponding Cohen’s $\delta$ effect sizes when comparing open and closed issues across different features. All features exhibit statistically significant differences ($p<0.01$). However, the effect sizes indicate the practical significance varies notably across features.
@@
  • has-label shows a large effect size ($\delta=0.94$), suggesting that closed issues/PRs are far more likely to be labeled compared to open issues and PRs.

### v2

@@ 3.2. Pre-processing Issues and Pull Requests
To enable reliable topic modeling of GitHub data, we perform a systematic preprocessing to clean and normalize the textual data. Specifically, we concatenate the title and body of each issue or pull request into a single document to preserve sufficient contextual information for topic extraction. We then remove code snippets using regular expressions, as such blocks often contain syntax elements irrelevant to semantic modeling. Similarly, we eliminate URLs, which typically link to external resources, and clean out residual HTML tags. We further normalize the corpus by removing special characters and punctuation. To address high-frequency but semantically uninformative words, we apply stopword removal using the NLTK stopword list (Hardeniya et al., 2016), and supplement it by filtering out project-specific terms such as nf-core and nextflow to prevent these frequently occurring system names from dominating the topic space. Finally, we apply lemmatization (Balakrishnan and Lloyd-Yemoh, 2014) using the spaCy en_core_web_sm model (Vasiliev, 2020) to convert words into their base forms, thus improving topic coherence and reducing lexical variation. This preprocessing ensures that the data retains its semantic richness while reducing noise, making it suitable for subsequent topic modeling and analysis. After preprocessing, we identify three samples with empty titles and bodies (these originally contained only URLs in the title and empty body). As they provide no meaningful information, we exclude them from the analysis. Consequently, the final dataset consists of 25,173 issues and pull requests.
@@ 3.3. Identifying Topics in nf-core Pipelines
Determining topic modeling parameters: In BERTopic, the nr_topics parameter allows users to control the number of topics by merging similar ones after their initial creation. However, prior work (Alam et al., 2025; Li et al., 2021) has emphasized the value of relying on clustering algorithms to produce a more natural topic structure. To enhance clustering quality, we first reduce the dimensionality of the document embeddings, as high-dimensional data can hinder performance. For this step, we employ UMAP (McInnes et al., 2018), which effectively preserves local semantic structure while projecting embeddings into a lower-dimensional space. We then apply HDBSCAN (McInnes et al., 2017), a density-based algorithm particularly well suited to UMAP projections. HDBSCAN is capable of identifying noise points and does not force every document into a cluster, making it especially effective for GitHub data.
For BERTopic modeling, we systematically tune UMAP and HDBSCAN parameters to balance topic granularity with semantic coherence. To evaluate the resulting topics, we employ the topic coherence score (Röder et al., 2015), a widely used metric that measures the semantic similarity among the top words in a topic. This measure helps distinguish between topics that are semantically interpretable and those that are merely artifacts of statistical inference. For UMAP, we vary n_neighbors (15–50) to balance local semantic similarity with a meaningful global structure, and adjust n_components (5–30) to retain sufficient semantic information while enabling efficient clustering. We use cosine distance, which aligns well with sentence-transformer embeddings. For HDBSCAN, we experiment with min_cluster_size values between 30 and 200, adopting Euclidean distance as it is well suited to UMAP-reduced spaces. To improve topic representation, we apply CountVectorizer with unigram and bigram models, consistent with prior studies (Alam et al., 2025; Li et al., 2021; Abdellatif et al., 2020).
We achieve the highest topic coherence score of 0.904 for our dataset using n_neighbors = 30, n_components = 15, and min_cluster_size = 150. With this configuration, BERTopic produces 13 distinct topics. However, BERTopic introduces a special -1 topic to aggregate documents that do not align well with any of the discovered topics. Such cases typically occur when documents contain excessive noise (e.g., stopwords or irrelevant content) or lack strong connections to the main topics the model extracts. Consequently, the -1 topic acts as a catch-all cluster for outlier documents. Since our dataset underwent extensive preprocessing to remove noisy and irrelevant elements prior to topic modeling, our main topic began with the label -1. The preprocessing and topic generation scripts are available in our replication package (Random, 2025). A detailed analysis of the resulting topics is presented in the Section 4.
The replication package is available in our online appendix (Random, 2025).
@@ 4. Results
In this section, we analyze the issues and pull requests from the nf-core pipelines. Our experiments are designed to address the following three research questions:
  • RQ1: What topics do users raise and discuss in the issues and pull requests of nf-core pipelines?
  • RQ2: How are issues and pull requests managed and addressed in nf-core pipelines, and how do different features influence their resolution?
  • RQ3: To what extent do the practitioners find the development activities relevant to the topics they discuss in the issues and pull requests challenging?
@@ 4.1. RQ1: What topics do users raise and discuss?
Motivation: Understanding the topics raised in issues and pull requests is essential to identifying the challenges and needs in nf-core pipeline development and maintenance. Since these pipelines are community-driven and widely used in bioinformatics, the issues and pull requests often reflect recurring technical difficulties, usability concerns, feature requests, and collaborative practices. Importantly, the development of nf-core pipelines differs substantially from that of traditional software systems: pipelines must integrate heterogeneous bioinformatics tools, manage complex data dependencies, and ensure reproducibility across diverse computational environments (e.g., containers, HPC clusters, cloud platforms). These requirements introduce domain-specific challenges, such as compatibility with reference datasets, versioned tools, and workflow execution engines, which are less prevalent in general-purpose software engineering (Di Tommaso et al., 2017; Arvanitou et al., 2021; Gruening et al., 2019). Therefore, the challenges developers and users face in scientific workflow repositories such as nf-core may differ from those encountered in traditional software projects. The motivation behind this RQ1 is to reveal the challenges, needs, and recurring concerns of the nf-core pipeline repositories.
Approach: To identify the key topics, we apply BERTopic modeling to our datasets, as outlined in Section 3. The modeling process yields 13 distinct topics along with the distribution of co-occurring words for each. We then manually assign meaningful, contextually accurate labels to the identified topics. Following established practices in prior work (Alam et al., 2025; Li et al., 2021; Openja et al., 2020; Yang et al., 2016; Bagherzadeh and Khatchadourian, 2019; Scoccia et al., 2021), the first author, having over five years of experience in nf-core pipelines and more than a decade of professional software development experience, proposes initial candidate labels. These are based on topic keywords and a manual review of at least 30 randomly selected representative issues and pull requests per topic. The proposed labels are subsequently refined through collaborative discussions with the second author, who has over nine years of experience in bioinformatics-related pipeline research, and an additional expert with more than two decades of expertise in empirical software engineering research. All disagreements were resolved through structured discussions. Through this iterative process, we finalize a set of labels that accurately capture the underlying semantics and contextual nuances of the identified topics.
@@
1. Pipeline Development and Integration: This is the most dominant topic users face when working with nf-core pipelines, focusing on the foundational tasks of initiating and integrating new workflows in the nf-core ecosystem. For example, in the hlatyping pipeline, an issue, nf-core pipelines should provide structured output information, proposed extending nf-core metadata to not only describe inputs but also provide structured output information, helping humans and automated systems anticipate outputs, and route results. Developers frequently scaffold pipelines using standardized templates, add files and test stubs, and ensure compliance with nf-core community guidelines. Keywords such as new tool, add test, pipeline, and documentation update highlight a strong emphasis on code quality and adherence to standards, with many issues and pull requests focusing on early development tasks.
2. Template Sync. and Automated Update Management: This is the second most frequent challenge, highlighting the need to keep pipelines synchronized with the latest nf-core template. Issues and pull requests in this topic involve propagating new conventions and updates, resolving merge conflicts, and ensuring community standards. For instance, the dualrnaseq pipeline issue titled as Pipeline needs updating to the latest template version required fixing the TEMPLATE branch and creating a sync pull request to resolve conflicts. Keywords such as conflict branch, make resolve, and automate attempt illustrate the hybrid nature of this process, where automation drives standardization but human intervention remains essential.
@@
5. Bug Fixing: Users often encounter challenges to ensure pipeline reliability by resolving bugs, adding tests, and maintaining representative datasets (e.g., fix bug, add test, nfcoretestdataset). For example, the pull request Bugfix MALT error retries addresses a memory-related issue by enabling the pipeline to retry when Java exits with code 1 due to heap space exhaustion, rather than aborting immediately.
6. Coordinating Contributions and Maintaining Pipeline Quality: Users often struggle to balance multiple tasks when contributing, such as updating documentation, fixing issues, and adding tests. Keywords like contribute, make branch, and appropriate delete reflect recurring coordination challenges across branches and the cleanup of outdated content. For example, the pull request Revert multiqc workaround restores original functionality by removing a temporary fix after a patch release resolved the underlying issue.
@@
10. Tool Development and Repository Maintenance: Users frequently encounter challenges when creating new tools and maintaining the repositories. Keywords such as make and new tool reflect to the complexity of initiating tool development, while add test, branch pipeline, and appropriate delete indicate difficulties in ensuring correctness, managing branches, and removing outdated files. For example, the pull request Add Kraken2 processes to build database and filter read illustrates these efforts.
@@
RQ1 Summary In RQ1, we identify 13 distinct topics through BERTopic modeling and manual validation. These topics encompass areas such as pipeline development and integration, template synchronization, debugging execution failures, tool and documentation maintenance, contribution coordination, genome data integration, module and parameter updates, subworkflow development, CI configuration, reporting/visualization, and version management. The taxonomy demonstrates that user discussions in nf-core repositories span the entire workflow lifecycle.
4.2. RQ2: Management and resolution of issues and pull requests in nf-core pipelines.
@@
Motivation: In nf-core pipelines, GitHub issues and pull requests serve as the primary channels for reporting bugs, requesting features, and contributing improvements. They reflect both the technical challenges users encounter and the collaborative processes used to resolve them. Understanding how these artifacts are managed is essential, as their resolution directly affects pipeline reliability, usability, and sustainability. As highlighted in RQ1, users raise issues and pull requests for diverse reasons, and ignoring these challenges can prevent users from successfully using these pipelines. While some cases, such as installing required packages, can be easily addressed, others, like tool development, CI management, and debugging, require coordinated effort and management.
While prior studies have explored issues and pull requests management in general open-source projects, pipeline repositories like nf-core pose unique, domain-specific challenges that may shape how issues are prioritized, resolved, or neglected. Investigating how these processes are managed and how features such as labeling, assignment, or code snippets influence resolution can provide important insights into the socio-technical dynamics of collaborative pipeline maintenance. Thus, RQ2 seeks to uncover the practices and strategies that sustain effective issue/pull request management in nf-core, offering lessons for both the nf-core community and the broader ecosystem of reproducible pipeline development.
@@
Figure 2. The distribution of (a) closed issues and pull requests rate, (b) ignored issues and pull requests rate, and (c) addressing time.
@@
How are the issues and pull requests addressed? We examine how issues and pull requests are managed in our dataset. Among 25,173 issues and pull requests, 22,500 (89.38%) are closed, and 2,673 (10.62%) remained open, with an overall average of 201.38 entries per repository. Figure 2(a) illustrates the distribution of the closed issues and pull requests rate (CIR) across repositories. Notably, 1.61% of repositories have achieved a CIR of 1.0, indicating that all reported issues and pull requests have been resolved. Moreover, repositories with a CIR exceeding 0.9 exhibit an average of 259.13 issues and pull requests, suggesting that repositories with higher closure rates also tend to manage a larger volume of entries. Overall, 48.39% of repositories demonstrate a CIR greater than 0.9, highlighting that nearly half of the repositories maintain a consistently high resolution rate.
@@
We find that 60.88% of closed issues and pull requests are self-closed, meaning that the individuals who reported them were able to resolve their concerns, often through subsequent discussion. In addition, we identify another type of case referred to as ignored issues and pull requests. These are cases that only the individuals who raised them are interested in, and receive no response from other users. These ignored cases account for 4.64 % of all issues and pull requests. We also define the ignored issues and pull requests rate as the ratio of ignored cases for each repository. This metric provides insight into the extent of active engagement by repository maintainers in addressing cases. The distribution of this rate across repositories is presented in Figure 2(b).
@@
We also analyze issues and pull requests addressing time, defined as the duration between creation and closure. Among 22,500 closed cases, the mean addressing time is 47.51 days with a standard deviation of 141.67 days, while the median is 2.89 days, indicating most are resolved quickly. However, a subset remains unresolved for extended periods. The maximum addressing time observed exceeds four years, highlighting significant variability in resolution dynamics. Figure 2(c) shows the distribution of the addressing time.
@@
To examine factors influencing issue and pull requests resolution time, we analyze two repository-level features following prior work (Yang et al., 2023): (1) the number of contributors and (2) the total number of issues and pull requests. We hypothesize that a larger contributor base may reduce addressing time. A negative Spearman correlation of –0.198 (p = 0.0283) between contributors and median closing time indicates that repositories with more contributors resolve issues slightly faster. Although this effect is statistically significant, the practical impact remains small. We further hypothesize that repositories with more issues and pull requests would exhibit longer addressing times, as a larger workload might burden maintainers. Contrary to this expectation, the results reveal a Spearman correlation of –0.218 (p = 0.0157) between the number of issues and pull requests and the median closing time. This negative correlation suggests that repositories handling a greater volume of issues/pull requests actually resolve them somewhat faster, though again, the effect size is modest.
@@
How are the issues and pull requests managed? GitHub offers several mechanisms to support maintainers in managing issues and pull requests. Following prior work (Yang et al., 2023), we focus on two commonly used strategies: (1) labeling, where issues and pull requests are tagged with labels, and (2) assigning, where issues and pull requests are allocated to specific individuals for resolution. In the remainder of this section, we use the terms labeling and assigning to refer to these two practices.
@@
We observe that 33.23% (8,366) of issues and pull requests are tagged with labels, while 17.13% are assigned to specific contributors. This indicates that a large portion of issues and pull requests do not make use of GitHub’s built-in management features. To further examine the use of labels, we note that GitHub provides nine default options: bug, documentation, duplicate, enhancement, help wanted, invalid, question, wontfix, and good first issue. Actively managed projects often go beyond these defaults by defining additional labels (e.g., DSL2). Across all pipeline repositories, we identify 154 distinct labels (including the defaults). We then calculate the frequency of each label, with the ranked distribution presented in Figure 3.
@@
Relationships between different features and the closure of issues and pull requests: We further examine how different characteristics of issues and pull requests relate to their likelihood of closure. Understanding these relationships offers practical insights for repository maintainers and contributors on improving the management and resolution of issues and pull requests. Specifically, we focus on the following features:
@@
1. has-label: Indicates whether an issue/pull request has a label. A labeled entry usually means that it has been noticed and read by the maintainers, thereby increasing the likelihood of being addressed.
@@
2. has-assignee: An assigned issue/pull request typically indicates clear responsibility, which may increase the likelihood of timely and successful resolution.
@@
3. title-length: Refers to the length of an issue or pull request title. Informative, longer titles provide more context, helping maintainers understand and resolve items more quickly.
@@
4. body-length: Detailed descriptions give maintainers essential context, aiding triage and resolution, and are often encouraged by project guidelines.
@@
5. has-code: Indicates if the issue or pull request includes code blocks, which help maintainers reproduce and resolve issues more effectively.
@@
6. has-URL: Indicates if the issue or pull request contains URLs, which provide helpful context through links to related issues, documentation, or external resources.
@@
7. num-contributors: Intuitively, repositories with a larger contributor base have greater collective capacity, which may increase the chances of timely issue/pull request resolution.
@@
Feature Measurement and Statistical Test: We compute the length of the issue/pull request title and body. Using regular expressions, we then identify whether an issue/pull request description contains code blocks or URLs. The number of contributors is obtained by querying the GitHub API to fetch metadata about the repository. Finally, we apply the Wilcoxon rank-sum test (Wilcoxon, 1945) to assess the statistical significance of the differences in feature distributions between open and closed issues and pull requests.
@@
Table 2. The significance of the differences between the feature values of open and closed issues and pull requests (Wilcoxon rank-sum test and Cohen’s effect sizes $\mathit{d}$).
Feature P-value Cohen’s $\mathit{d}$
@@
We also estimate Cohen’s effect size ($\mathit{d}$) to quantify the magnitude of differences between open and closed issues and pull requests. Following established guidelines, values of $\mathit{d}<0.2$ are regarded as negligible, $0.2\leq|\mathit{d}|<0.5$ as small, $0.5\leq|\mathit{d}|<0.8$ as medium, and $|\mathit{d}|\geq 0.8$ as large. Table 2 reports the results of the Wilcoxon rank-sum test and corresponding Cohen’s $\mathit{d}$ effect sizes when comparing open and closed issues and pull requests across different features. All features exhibit statistically significant differences ($p<0.01$). However, the effect sizes indicate that the practical significance varies notably across features.
@@
  • has-label shows a large effect size ($\mathit{d}=0.94$), suggesting that closed issues/pull requests are far more likely to be labeled compared to open issues and pull requests.